Determine the given and compute the appropriate test statistic of the problem below. Construct the rejection region of the problem below


A manufacturer of face mask has developed a new face mask design. He claims that the new design has an average profit increase of 10% with a standard deviation of 3%. Test the hypothesis that the new face mask design average profit increase of is not 10% if a random sample of 50 face mask is tested with an average profit increase of 4%. Use 10% level of significance.


1
Expert's answer
2022-05-10T07:01:20-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=10"

"H_1:\mu\not=10"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.1," and the critical value for a two-tailed test is "z_c = 1.6449."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{4-10}{3/\sqrt{50}}=-14.142"

Since it is observed that "|z| = 14.142 > 1.6449=z_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=2P(Z<-14.142)=0,"  p and since "p = 0 < 0.1=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 10, at the "\alpha = 0.1" significance level.